f(x)=f(x+1)/3 + 5/f(x+2) lim x tends to infinity how to solve it???

Grade:12

1 Answers

Sher Mohammad IIT Delhi
9 years ago
let limit f(x)=k, k=k/3+ 5/k
3k^2=k^2+15
k^2=15
k=+-sqrt(15)

sher mohammad
askiitians faculty
b.tech, IIT Delhi